1/2

Heinrich Heine,

The Lorelei

translated by Mark Twain

An ancient legend of the Rhine

I cannot divine what it meaneth, This haunting nameless pain: A tale of the bygone ages Keeps brooding through my brain:

The faint air cools in the gloaming, And peaceful flows the Rhine, The thirsty summits are drinking The sunset’s flooding wine;

The loveliest maiden is sitting High-throned in yon blue air, Her golden jewels are shining, She combs her golden hair;

She combs with comb that is golden, And sings a weird refrain That steeps in a deadly enchantment The listener’s ravished brain:

The doomed in his drifting shallop, Is tranced with the sad sweet tone, He sees not the yawing breakers, He sees but the maid alone:

…

2/2

The pitiless billwos engulf him!- So perish sailor and bark; And this, with her baleful singing, Is the Lorelei’s gruesome work.

Translation of the German poem “Die Lorelei” by Heinrich Heine (∗1797-12-13 – †1856-02-17) into English by Mark Twain.

Arg-2-11 (2003-10-13 04:42:59)

A Tramp Abroad. Vol 1-2. Leibzig: Tauchnitz, 1880 Band I, Mark Twain 1880 (rf. http: // www. loreley. com/ loreley/ marctwai. htm )

http://www.poezio.net/version?poem-id=2&version-id=11